What is the vertex given the equation: y=2/3 square root of x+4-2

Vertex:

1
Expert's answer
2021-03-12T12:21:59-0500



our function is

"y=\tfrac{2}{3}\sqrt{\smash[b]{x+4}}-2"

Standard form of the square root function is

"y=a\sqrt{\smash[b]{b(x-h)}}+k" , where

"b=\u00b11" and

"h, k" - vertex

Convert our function

"y=\tfrac{2}{3}\sqrt{\smash[b]{x+4}}-2=\tfrac{2}{3}\sqrt{\smash[b]{-(-x-4)}}-2"

Therefore vertex is "(-4, -2)"

Also we can see vertex on the graph of the function.